An aquarium holding African cichlids is 4 inches higher than it is wide. It's length is 29 inches, and it's volume is 4060 in.³
Found 2 solutions by algebrahouse.com, Alan3354:
Answer by algebrahouse.com(1659)

w = width
w + 4 = height
29 = length

Volume of an aquarium (rectangular prism)
V = lwh

29(w)(w + 4) = 4060 {substituted into volume formula}
w(w + 4) = 140 {divided each side by 29}
w² + 4w = 140 {used distributive property}
w² + 4w - 140 = 0 {subtracted 140 from each side}
(w + 14)(w - 10) = 0 {factored into two binomials}
w + 14 = 0 or w - 10 = 0 {set each factor equal to 0}
w = -14 or w = 10 {solved each equation for w}
w = 10 only {width of an aquarium cannot be negative}
w + 4 = 14 {substituted 10, in for w, into w + 4}

width = 10 inches
height = 14 inches
length = 29 inches

For more help from me, visit: www.algebrahouse.com



Answer by Alan3354(69443)

it's = it is